Good evening,
My question really involves function operations and I was wondering if anybody could help me. The question is suppose that f(x)=x^2-x and g(x)=2x+1. What is f times g(x). The answer I received was 2x^3+x^2+x am I correct. Please get back to me when you can. Thank you in advance.

Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39625)

Try again and watch the signs carefully.


In a grid or lattice form,


_______________x^2__________-x
-
-
____2x________2x^3_________-2x^2
-
-
____1_________x^2__________-x

Answer by stanbon(75887)

suppose that f(x)=x^2-x and g(x)=2x+1. What is f times g(x). The answer I received was 2x^3+x^2+x am I correct.
-------
f*g(x) = (x^2-x)(2x+1) = x^2(2x+1) -x(2x+1)
----
= 2x^3 + x^2 - 2x^2 -x
----
= 2x^3 - x^2 - x
======================
Cheers,
Stan H.
---------------------